OpenStudy (anonymous):

A diesel engine works at a high compression ratio to compress air until it reaches a temperature high enough to ignite the diesel fuel. Suppose the compression ratio (ratio of volumes) of a specific diesel engine is 24 to 1. If air enters a cylinder at 1 atm and is compressed adiabatically, the compressed air reaches a pressure of 66.0 atm. Assuming that the air enters the engine at room temperature (21.9°C) and that the air can be treated as an ideal gas, find the temperature (in K) of the compressed air.

OpenStudy (anonymous):

\[P1^(1-\gamma)*T1^\gamma=P2^(1-\gamma)*T2^\gamma\] P1=1atm;T1=(273+21.9);P2=66atm; now get T2